OpenStudy (anonymous):

The sum of the angle measures of any triangle is 180. the second angle measures three times first, and the 3rd angle measures 30 less than the second. what is the measure of each of the angle?

OpenStudy (anonymous):

Do the same thing i showed you in your last question

OpenStudy (anonymous):

solve the equations : b=3a c=b/30 a+b+c = 180

OpenStudy (anonymous):

1st=x 2nd=3x 3rd=3x-30 x+3x+3x-30=180 7x=210 x=30 1st=30 2nd=90 3rd=60

OpenStudy (anonymous):

thanks